# PET2Rep: Towards Vision-Language Model-Drived Automated Radiology Report Generation for Positron Emission Tomography

### v1

@@ Abstract
Positron emission tomography (PET) is a cornerstone of modern oncologic and neurologic imaging, distinguished by its unique ability to illuminate dynamic metabolic processes that transcend the anatomical focus of traditional imaging technologies. Radiology reports are essential for clinical decision making, yet their manual creation is labor-intensive and time-consuming. Recent advancements of vision-language models (VLMs) have shown strong potential in medical applications, presenting a promising avenue for automating report generation. However, existing applications of VLMs in the medical domain have predominantly focused on structural imaging modalities, while the unique characteristics of molecular PET imaging have largely been overlooked. To bridge the gap, we introduce PET2Rep, a large-scale comprehensive benchmark for evaluation of general and medical VLMs for radiology report generation for PET images. PET2Rep stands out as the first dedicated dataset for PET report generation with metabolic information, uniquely capturing whole-body image-report pairs that cover dozens of organs to fill the critical gap in existing benchmarks and mirror real-world clinical comprehensiveness. In addition to widely recognized natural language generation metrics, we introduce a series of clinical efficiency metrics to evaluate the quality of radiotracer uptake pattern description in key organs in generated reports. We conduct a head-to-head comparison of 30 cutting-edge general-purpose and medical-specialized VLMs. The results show that the current state-of-the-art VLMs perform poorly on PET report generation task, falling considerably short of fulfilling practical needs. Moreover, we identify several key insufficiency that need to be addressed to advance the development in medical applications. We believe PET2Rep will serve as a platform for the development and application of VLMs for PET imaging, accelerating the development of trustworthy reporting tools that can genuinely alleviate radiologist burden and enhance patient care. Project page: https://github.com/YichiZhang98/PET2Rep.
@@ Introduction
The recent surge in vision-language models (VLMs) has attracted interest from the medical community, where these models hold great potential to transform various aspects of clinical practice like automatic medical report generation (Zhang et al. 2024). Leveraging the power of large-scale pretraining, VLMs can analyze medical images and generate corresponding textual descriptions, effectively bridging the gap between visual data and clinical language. However, existing applications of VLMs in the medical domain have predominantly focused on structural imaging modalities (Liu et al. 2024; Hamamci et al. 2024; Zhu et al. 2025b), while the unique characteristics and clinical value of PET imaging have largely been overlooked in the current research landscape. The analysis of PET images poses unique challenges due to the need to integrate functional and anatomical information and specialized knowledge required to interpret tracer uptake patterns (Coleman et al. 2010). Despite its clinical significance, there is a dearth of research exploring the application of VLMs in PET imaging, leaving a significant gap in the development of automated reporting systems for this critical modality. Addressing this gap could potentially streamline the PET reporting process, enhance diagnostic accuracy, and improve patient outcomes by providing timely and comprehensive reports (Matsubara et al. 2022).
@@
To address these challenges, we introduce PET2Rep, a large-scale comprehensive benchmark for the evaluation of radiology report generation for PET imaging. The evaluation is conducted on a large-scale multimodal dataset of 565 cases with paired PET/CT imaging and corresponding radiology reports. Compared with existing medical benchmarks, the key advantages of PET2Rep can be concluded in the following three aspects.
@@
The First PET Report Dataset. PET2Rep is the first dataset dedicated to PET report generation. Unlike other modalities such as X-ray and CT which primarily focus on anatomical structures and volumetric measurements, PET operates at the molecular level, enabling the assessment of metabolic information. This unique feature allows for early disease identification, often before anatomical changes are visible on other imaging modalities (Gatidis et al. 2024). Given the high cost of PET scans and the need for specialized expertise in report writing, there is currently no relevant dataset available, which highlights the importance of PET2Rep in advancing research in this field.
@@
Whole-Body Imaging with Radiology Reports. Existing medical imaging benchmarks are often limited to specific anatomical domains. For instance, chest X-ray report generation primarily address thoracic pathologies (Liu et al. 2024), while those for CT reports concentrate on the volume and morphology of organs and lesions in the chest (Hamamci et al. 2024) or abdominal regions (Bassi et al. 2025). In contrast, PET2Rep encompasses a much broader anatomical scope, with images ranging from the head and neck to the proximal limbs. Consequently, its corresponding reports provide detailed evaluations of dozens of organs body-wide, demanding a more extensive scope of medical knowledge for accurate interpretation. This holistic approach more closely simulates real-world oncology practice, where radiologists conduct comprehensive patient assessments rather than focusing on isolated anatomical areas.
@@
Data Collection from Real Clinical Scenarios. PET2Rep is collected from real clinical scenarios to bridge the gap between algorithmic benchmarks and actual clinical practice. Many existing medical multimodal benchmarks are developed from public imaging archives (Sepehri et al. 2024; Chen et al. 2024). These frameworks often generate tasks that probe for superficial understanding of the image, such as identifying the imaging modality or naming marked organs, rather than complex clinical reasoning (Ye et al. 2024; Zhou et al. 2025). Such scenarios test for basic medical knowledge and differ significantly from the complex demands of a real clinical workflow. In contrast, PET2Rep incorporates data directly from the clinical setting, ensuring that the benchmark authentically reflects the challenges radiologists encounter in their daily work. This ensures the authenticity and clinical relevance of the PET2Rep benchmark while minimizing the risk of data leakage, thereby reflecting the generalization performance of VLMs in real-world clinical scenarios.
@@
To make a comprehensive evaluation of the performance of VLMs, we establish a standardized evaluation pipeline for PET radiology report generation. We formulate a prompting framework incorporating essential elements including imaging modality specifications and clinical objectives and design a structured report template aligned with radiological training protocols. This approach ensures faithful translation of image-derived information into formatted reports that maintain consistency with expert-generated radiological reports. We conduct a comprehensive evaluation state-of-the-art models, including 19 general purpose and 11 medical-specific VLMs on PET2Rep benchmark. The experimental results show that current cutting-edge VLMs exhibit suboptimal performance on the task, falling considerably short of fulfilling real-world requirements. Furthermore, our analysis reveals several critical limitations that must be tackled to drive progress in clinical applications.
@@
Figure 2: Pipeline of the PET2Rep benchmark for evaluation of VLM-based PET radiology report generation. First, PET/CT images are analyzed by VLMs with a designed prompt format to include necessary information such as image modality, clinical task, and designed report template based on radiologist training guidelines. Then the generated reports are evaluated against the ground-truth reports with widely recognized natural language generation (NLG) metrics and a novel clinical efficiency (CE) metric for PET imaging. We further conduct manual scoring by radiologists for more comprehensive evaluation.
@@
Positron Emission Tomography (PET) is a clinical imaging technique that reveals ongoing metabolic processes in the body by detecting gamma photons generated from positron annihilation after injecting radioactive tracers. As the most widely used tracer, fluorodeoxyglucose (FDG) assesses local glucose uptake to evaluate organ metabolism and detect tumor metastasis, enabling monitoring of treatment progress (Ren et al. 2019). Clinically, PET is primarily used for early tumor screening for cancer detection (Gatidis et al. 2024; Peng et al. 2023), organ metabolic function assessment (Xue et al. 2024; Zhang et al. 2025), and treatment monitoring (van der Geest et al. 2021). The unique metabolic imaging capability of PET provides indispensable insights for disease diagnosis and treatment optimization in clinical applications.
@@
Vision-Language Models (VLMs) have emerged as a transformative force in artificial intelligence, adeptly bridging the gap between visual perception and natural language understanding (Zhang et al. 2024). The swift progress in VLM development is largely attributed to innovative pre-training strategies and architectural designs, which have demonstrated remarkable capabilities across a wide array of tasks such as visual question answering and image captioning (Chen et al. 2022; Feng et al. 2025; Lin et al. 2025). Beyond general-purpose vision tasks, VLMs are making significant inroads into specialized fields like medical image analysis (Peng et al. 2025). VLMs can generate diagnostic reports, answer clinical questions, and highlight regions of interest, offering substantial support to healthcare professionals and promising to enhance the efficiency and accuracy of medical diagnoses (Zhang, Shen, and Jiao 2024; Jiang et al. 2024; Lin, Xu, and Qin 2025).
@@
Overview
In this work, we introduce PET2Rep, the first specialized benchmark designed to evaluate the performance of VLMs for generating radiology reports from PET images. PET2Rep is the first PET dataset with paired structured radiology reports. A key distinction of PET2Rep is its data sourcing. Unlike benchmarks that rely on data aggregated from online repositories, all data in our work were meticulously collected from actual clinical cases across multiple medical centers. This approach guarantees the authenticity and clinical relevance of the benchmark. Furthermore, by sourcing directly from clinical settings, we mitigate the risk of data leakage, ensuring that PET2Rep accurately reflects the complexity and diversity of real-world radiological practice. The comprehensive workflow of the PET2Rep benchmark is detailed in Table. 2 and will be elaborated upon in the subsequent sections.
@@
We collected 565 cases of 3D whole-body FDG PET/CT imaging from one local medical center, which is the most widely used PET tracer in oncology. As a non-specific tracer, FDG can be used for whole-body imaging to reflect tissue glucose metabolism, which makes the imaging useful in assessing the systemic distribution and metastasis of tumors. Structured radiology reports are constructed using a radiologist-designed template drafted, which plays a crucial role in assisting physicians to analyze whole-body PET/CT imaging in a standardized format, so as to enhance clarity and aid in medical decisions. The report presents a detailed, objective account of the observations made from the whole-body PET/CT images in a systematic, head-to-toe order, ensuring no anatomical region is overlooked. It lists all the detected abnormalities without offering any interpretative conclusions. This part is factual and serves as the raw data foundation for further assessment. More detailed information and examples of the dataset are shown in the Appendix.
@@
To ensure the accuracy and reliability of multi-modal image analysis, rigorous pre-processing of the data is essential to bring the data into a consistent and interpretable format. In PET/CT imaging, a key pre-processing step involves resampling the CT images to match the lower spatial resolution of the PET images. This coregistration process ensures that both imaging modalities share the same matrix size and that the voxels in each dataset correspond to the same anatomical location. After resampling, the intensity of the CT images is standardized using z-score. Another vital preprocessing step is the normalization of the PET data. The raw PET images, which measure radioactivity in counts, are converted to Standardized Uptake Values (SUV), which is a widely accepted quantitative metric in PET imaging that adjusts the measured radioactivity concentration for factors such as the injected radiotracer dose and the patient’s body weight (Lucignani, Paganelli, and Bombardieri 2004).
To emulate the clinical diagnostic workflow, we implement a process to fuse PET and CT scans. This technique combines the functional information from PET with the anatomical detail from CT, mirroring how radiologists interpret these images for diagnosis. This process yields a single composite view for visualizing functional information within its precise anatomical context. This integration is critical for accurately localizing areas of abnormal radiotracer uptake and enables a more holistic and integrated assessment of the patient’s condition.
@@
Given that the original PET/CT images are three-dimensional, while most existing VLMs are designed for 2D images. To bridge the gap, it becomes necessary to select representative 2D slices from the 3D medical imaging. We chose the coronal plane as the view for slice sampling, a decision grounded in clinical practice where radiologists frequently use this view for a comprehensive, head-to-toe assessment. The coronal plane is ideal for this purpose because it provides a complete overview of the entire body and the majority of key organs. By analyzing multiple coronal slices, our model can effectively evaluate the critical information contained within the full 3D scan. In this work, we simulate the diagnostic process of radiologists and design two strategies for report generation as follows.
Input Separate PET and CT Images.
In this strategy, we maintain the distinction between functional and anatomical information by providing the model with two distinct, parallel inputs. For each anatomical location of interest, we extract a corresponding pair of 2D slices: a grayscale slice from the CT volume for structural context, and a pseudo-colored slice from the PET volume to highlight metabolic activity. Specifically, we identify three key locations for analysis, resulting in a total input of six images for the VLMs (a PET/CT pair for each location). This dual-input approach compels the model to learn the complex correlations between anatomical structure and functional uptake, mirroring the cognitive process of a radiologist integrating two different sets of images.
Input Fused PET/CT Images.
This strategy involves pre-integrating the multimodal information into a single image before presenting it to the model. For each selected location, we generate a fused image by superimposing the pseudo-colored PET slice directly onto its corresponding grayscale CT slice. In our implementation, we create these fused images for three key locations, providing the model with a total of three input PET/CT images. Each resulting image offers an immediate, composite view where metabolic hotspots are displayed within their precise anatomical context. This method simulates the final visualization that radiologists use for diagnosis. By providing a single, pre-fused input for each location, we reduce the model’s burden of performing the fusion itself, allowing it to focus directly on interpreting the combined functional and structural patterns.
@@
In this study, we evaluated a range of VLMs for PET report generation, encompassing both medical-specific models and general-purpose models including open-source options and those accessible via proprietary APIs. The weights of open-source models were sourced from respective official Hugging Face repositories. To ensure that VLMs can understand the task and output reports in a standardized format like radiologists, we formulate a prompt format for the benchmark. This format includes necessary information such as imaging modality and clinical tasks, as well as a report template designed based on radiologist training guidelines. This design ensures that the interpretation of image information can be reasonably converted into a format consistent with manually annotated radiological reports. Our evaluation of VLMs was conducted in a zero-shot setting, which serves as a stringent test of the models’ generalization ability, revealing how well they can handle complex medical imaging tasks without any task-specific fine-tuning. All tests were executed using NVIDIA A100 GPUs with 80GB of memory to ensure consistency, laying a solid foundation for accurate and efficient evaluation of the models’ performance.
@@
General-Purpose VLMs.
General-purpose VLMs are designed for broad applicability across a wide range of multimodal tasks. These models are characterized by their large-scale architectures and training on extensive, diverse datasets, enabling them to excel at visual understanding and reasoning tasks. Their primary strength lies in their versatility and scalability. We select following general-purpose VLMs for evaluation, including Qwen2.5-VL series (Bai et al. 2025), InternVL3 series (Zhu et al. 2025a), Yi-VL series (Young et al. 2024), LLaVA-V1.5 (Liu et al. 2023a), OmniLMM (Yu et al. 2024), VisualGLM (Du et al. 2022) and Deepseek-VL2(Wu et al. 2024) .
Medical-Specific VLMs.
In contrast to general-purpose models, medical-specific VLMs are meticulously tailored for the complexities and high-stakes nature of clinical applications. These models prioritize domain adaptation, clinical context, and the integration of specialized medical knowledge. Their development involves training on curated medical datasets and aligning their outputs with diagnostic workflows to ensure accuracy and reliability in healthcare scenarios. We select following medical-specific VLMs for evaluation, including LLaVA-Med (Li et al. 2023), Med-Flamingo (Moor et al. 2023), Qilin-Med-VL (Liu et al. 2023b), RadFM (Wu et al. 2023), MedDr (He et al. 2024), HuatuoGPT-Vision (Chen et al. 2024), MedVLM-R1 (Pan et al. 2025), and latest MedGemma series (Sellergren et al. 2025) and Lingshu series (Xu et al. 2025).
Closed-Source Commercial VLMs.
Closed-source commercial VLMs are developed and maintained by enterprises, with their source codes not publicly available. These models are made accessible to users primarily through API interfaces, allowing for convenient integration into various applications. We select following closed-source VLMs for evaluation, including Gemini 2.5 Pro (Comanici et al. 2025), GPT-4o (Hurst et al. 2024), Moonshot-v1 (MoonshotAI 2025) and Qwen-VL-Max (Bai et al. 2023) for evaluation.

### v2

@@ Abstract
Positron emission tomography (PET) is a cornerstone of modern oncologic and neurologic imaging, distinguished by its unique ability to illuminate dynamic metabolic processes that transcend the anatomical focus of traditional imaging technologies. Radiology reports are essential for clinical decision making, yet their manual creation is labor-intensive and time-consuming. Recent advancements of vision-language models (VLMs) have shown strong potential in medical applications, presenting a promising avenue for automating report generation. However, existing applications of VLMs in the medical domain have predominantly focused on structural imaging modalities, while the unique characteristics of molecular PET imaging have largely been overlooked. To bridge the gap, we introduce PET2Rep, a large-scale comprehensive benchmark for evaluation of general and medical VLMs for radiology report generation for PET images. PET2Rep stands out as the first dedicated dataset for PET report generation with metabolic information, uniquely capturing whole-body image-report pairs that cover dozens of organs to fill the critical gap in existing benchmarks and mirror real-world clinical comprehensiveness. In addition to widely recognized natural language generation metrics, we introduce a series of clinical efficacy metrics to evaluate the quality of radiotracer uptake pattern description in key organs in generated reports. We conduct a head-to-head comparison of 30 cutting-edge general-purpose and medical-specialized VLMs. The results show that the current state-of-the-art VLMs perform poorly on PET report generation task, falling considerably short of fulfilling practical needs. Moreover, we identify several key insufficiency that need to be addressed to advance the development in medical applications. We believe PET2Rep will serve as a platform for the development and application of VLMs for PET imaging, accelerating the development of trustworthy reporting tools that can genuinely alleviate radiologist burden and enhance patient care. Project page: https://github.com/YichiZhang98/PET2Rep.
@@ Introduction
The recent surge in vision-language models (VLMs) has attracted interest from the medical community, where these models hold great potential to transform various aspects of clinical practice like automatic medical report generation (Zhang et al. 2024). Leveraging the power of large-scale pretraining, VLMs can analyze medical images and generate corresponding textual descriptions, effectively bridging the gap between visual data and clinical language. However, existing applications of VLMs in the medical domain have predominantly focused on structural imaging modalities (Liu et al. 2024; Hamamci et al. 2024; Zhu et al. 2025b), while the unique characteristics and clinical value of PET imaging have largely been overlooked in the current research landscape. As the analysis of PET images poses unique challenges due to the need to integrate functional and anatomical information and specialized knowledge required to interpret tracer uptake patterns (Coleman et al. 2010; Matsubara et al. 2022), it is worth rethinking that How Far are VLMs from Effective Radiology Report Generation for Positron Emission Tomography Imaging?
@@
To answer this question, we introduce PET2Rep, a comprehensive benchmark for the evaluation of radiology report generation for PET imaging. Compared with existing medical benchmarks, the key advantages of PET2Rep can be concluded in the following three aspects.
@@
The First PET/CT Report Dataset. PET2Rep is the first dataset dedicated to PET/CT report generation. Unlike other modalities like X-ray and CT which primarily focus on anatomical structures, PET operates at the molecular level, enabling the assessment of metabolic information. This unique feature allows for early disease identification, often before anatomical changes are visible on other imaging modalities (Gatidis et al. 2024). PET2Rep is a large-scale multi-modal dataset of 565 cases with paired PET, CT and corresponding radiology reports. Given the high cost of PET/CT scans and the need for specialized expertise in report writing, there is currently no relevant dataset available, which highlights the importance of PET2Rep in advancing research in this field.
@@
Whole-Body Imaging with Radiology Reports. Existing medical imaging benchmarks are often limited to specific anatomical domains. For instance, chest X-ray report generation primarily address thoracic pathologies (Liu et al. 2024), while those for CT reports concentrate on the volume and morphology of organs and lesions in the chest (Hamamci et al. 2024) or abdominal regions (Bassi et al. 2025). In contrast, PET2Rep encompasses a much broader anatomical scope, with images ranging from the head and neck to the proximal limbs. Consequently, its corresponding reports provide detailed evaluations of dozens of organs body-wide, demanding a more extensive scope of medical knowledge for accurate interpretation, as shown in Figure.1. This holistic approach more closely simulates real-world oncology practice, where radiologists conduct comprehensive assessments rather than focusing on isolated areas.
@@
Data Collection from Clinical Scenarios. Many existing medical multimodal benchmarks are developed from public imaging archives (Sepehri et al. 2024; Chen et al. 2024). These frameworks often generate tasks that probe for superficial understanding of the image, such as identifying the imaging modality or naming marked organs, rather than complex clinical reasoning (Ye et al. 2024; Zhou et al. 2025). Such scenarios test for basic medical knowledge and differ significantly from the complex demands of a real clinical workflow. In contrast, PET2Rep is collected from real clinical scenarios and incorporates data directly from the clinical setting, ensuring that the benchmark authentically reflects the challenges radiologists encounter in their daily work. This ensures the authenticity and clinical relevance of the PET2Rep benchmark while minimizing the risk of data leakage, thereby reflecting the generalization performance of VLMs in real-world clinical scenarios.
@@
To make a comprehensive evaluation of the performance of VLMs, we establish a standardized evaluation pipeline for PET/CT radiology report generation. We formulate a prompting framework incorporating essential elements including imaging modality specifications and clinical objectives and design a structured report template aligned with radiological training protocols. This approach ensures faithful translation of image-derived information into formatted reports that maintain consistency with expert-generated radiological reports. We conduct a comprehensive evaluation state-of-the-art models, including 19 general purpose and 11 medical-specific VLMs on PET2Rep benchmark. The experimental results show that current cutting-edge VLMs exhibit suboptimal performance on the task, falling considerably short of fulfilling real-world requirements. Furthermore, our analysis reveals several critical limitations that must be tackled to drive progress in clinical applications.
@@
Figure 2: Pipeline of the PET2Rep benchmark for evaluation of VLM-based PET radiology report generation. First, PET/CT images are analyzed by VLMs with a designed prompt format to include necessary information such as image modality, clinical task, and designed report template based on radiologist training guidelines. Then the generated reports are evaluated against the ground-truth reports with widely recognized natural language generation (NLG) metrics and a novel clinical efficacy (CE) metric for PET imaging. We further conduct manual scoring by radiologists for more comprehensive evaluation.
@@
Positron Emission Tomography (PET) is a clinical imaging technique that reveals ongoing metabolic processes in the body by detecting gamma photons generated from positron annihilation after injecting radioactive tracers. As the most widely used tracer, fluorodeoxyglucose (FDG) assesses local glucose uptake to evaluate organ metabolism and detect tumor metastasis, enabling monitoring of treatment progress (Ren et al. 2019). Clinically, PET is primarily used for early tumor screening for cancer detection (Gatidis et al. 2024; Peng et al. 2023), organ metabolic function assessment (Xue et al. 2024; Zhang et al. 2025), and treatment monitoring (van der Geest et al. 2021). The unique metabolic imaging capability provides indispensable insights for disease diagnosis and treatment optimization in clinical applications.
@@
Vision-Language Models (VLMs) have emerged as a transformative development in artificial intelligence, effectively bridging the gap between visual perception and natural language understanding (Zhang et al. 2024). The swift progress in VLM development is largely attributed to innovative pre-training strategies and architectural designs, which have demonstrated remarkable capabilities across a wide array of tasks such as visual question answering and image captioning (Chen et al. 2022; Feng et al. 2025; Lin et al. 2025). Beyond general-purpose vision tasks, VLMs are making significant inroads into specialized fields like medical image analysis (Peng et al. 2025). VLMs can generate diagnostic reports, answer clinical questions, and highlight regions of interest, offering substantial support to healthcare professionals and promising to enhance the efficacy and accuracy of medical diagnoses (Zhang, Shen, and Jiao 2024; Jiang et al. 2024; Lin, Xu, and Qin 2025).
@@
We introduce PET2Rep, a comprehensive benchmark designed to evaluate the performance of VLMs for generating radiology reports from PET images. PET2Rep is the first PET/CT dataset with paired structured radiology reports. A key distinction of PET2Rep is its data sourcing. Unlike benchmarks that rely on data aggregated from online repositories, all data in our work were meticulously collected from actual clinical scenarios. This approach guarantees the authenticity and clinical relevance of the benchmark. Furthermore, by sourcing directly from clinical settings, we mitigate the risk of data leakage, ensuring that PET2Rep accurately reflects the complexity and diversity of real-world radiological practice. The comprehensive workflow of the PET2Rep benchmark is detailed in Figure. 2 and will be elaborated upon in the subsequent sections.
@@
We collect 565 cases of 3D whole-body FDG PET/CT imaging from one local medical center, which is the most widely used PET tracer in oncology. As a non-specific tracer, FDG can be used for whole-body imaging to reflect tissue glucose metabolism, which makes the imaging useful in assessing the systemic distribution and metastasis of tumors. Structured radiology reports are constructed based on radiologist-designed templates, which play a pivotal role in assisting physicians to interpret whole-body PET/CT scans in a standardized and organized manner, thereby enhancing clarity and supporting clinical decision-making. The report provides a detailed and objective description of the findings from the PET/CT images in a systematic, head-to-toe sequence, ensuring that no anatomical region is overlooked. It enumerates all detected abnormalities without offering interpretative conclusions, serving as a factual foundation for subsequent clinical assessment. More detailed information and examples of the dataset are shown in the Appendix.
@@
To ensure the accuracy and reliability of multi-modal image analysis, rigorous data pre-processing is indispensable for bringing all imaging modalities into a consistent and interpretable format. In PET/CT imaging, a critical pre-processing step involves resampling the CT images to match the lower spatial resolution of the PET images. This coregistration process aligns both modalities to a common matrix size, ensuring voxel-wise correspondence across datasets. Following resampling, the CT intensities are standardized using z-score normalization to reduce inter-scan variability. Additionally, normalization of PET data is performed by converting the raw radioactivity counts into Standardized Uptake Values (SUV) a widely adopted quantitative metric in PET imaging that accounts for factors such as the injected radiotracer dose and the patient’s body weight (Lucignani, Paganelli, and Bombardieri 2004). To emulate the clinical diagnostic workflow, we implement a fusion process that integrates PET and CT scans. This approach combines the functional information from PET with the anatomical detail provided by CT, reflecting the manner in which radiologists interpret these modalities in clinical practice. The resulting composite image enables visualization of metabolic activity within its precise anatomical context. Such integration is essential for accurately localizing regions of abnormal radiotracer uptake and facilitating a comprehensive assessment of the patient’s condition.
@@
Given that the original PET/CT images are three-dimensional, while most existing VLMs are designed for 2D images, it becomes necessary to select out representative 2D slices from the 3D imaging. In this study, we select the coronal plane as the view for slice sampling, following clinical conventions in which radiologists commonly utilize this view for comprehensive head-to-toe assessments. The coronal plane offers an optimal perspective, capturing the global anatomical context and encompassing the majority of key organs. By analyzing multiple coronal slices, VLMs can effectively capture the salient information embedded within the full 3D scan. Building upon this design, we further emulate the diagnostic process of radiologists and design two strategies for report generation as described below.
Input Separate PET and CT Images. In this strategy, we maintain the distinction between functional and anatomical information by providing the model with two distinct, parallel inputs. For each anatomical location of interest, we extract a corresponding pair of 2D slices: a grayscale slice from the CT volume for structural context, and a pseudo-colored slice from the PET volume to highlight metabolic activity. Specifically, we identify three key locations for analysis, resulting in a total input of six images for the VLMs (a PET/CT pair for each location). This dual-input approach compels the model to learn the complex correlations between anatomical structure and functional uptake, mirroring the cognitive process of a radiologist integrating two different sets of images.
Input Fused PET/CT Images. This strategy involves pre-integrating the multimodal information into a single image before presenting it to the model. For each selected location, we generate a fused image by superimposing the pseudo-colored PET slice directly onto its corresponding grayscale CT slice. In our implementation, we create these fused images for three key locations, providing the model with a total of three input PET/CT images. Each fused image presents an immediate composite view, in which metabolic hotspots are precisely localized within their anatomical context. This method simulates the final visualization that radiologists use for diagnosis. By supplying the model with pre-fused inputs, we eliminate the need for it to learn the fusion process, enabling it to focus directly on understanding the integrated functional and structural patterns.
@@
In this study, we evaluate a range of VLMs encompassing both medical-specific and general-purpose models including open-source options and those accessible via proprietary APIs. The weights of open-source models were sourced from respective official Hugging Face repositories. To guide the models in generating radiology-style reports, we design a standardized prompt format specifying the imaging modality, clinical task, and a report template derived from radiologist training guidelines. This ensures that image interpretations are expressed in a format consistent with manually authored radiological reports. Our evaluation are conducted in a zero-shot setting, which serves as a stringent test of the models’ generalization ability, revealing how well they can handle complex medical imaging tasks without any task-specific fine-tuning. All tests were executed using NVIDIA A100 GPUs with 80GB of memory.
@@
General-Purpose VLMs. General-purpose VLMs are trained on large and diverse datasets to perform a wide spectrum of multimodal tasks. Their broad training enables strong visual understanding and reasoning capabilities, with versatility and scalability. We select following general-purpose VLMs for evaluation, including Qwen2.5-VL series (Bai et al. 2025), InternVL3 series (Zhu et al. 2025a), Yi-VL series (Young et al. 2024), LLaVA-V1.5 (Liu et al. 2023a), OmniLMM (Yu et al. 2024), VisualGLM (Du et al. 2022) and Deepseek-VL2 (Wu et al. 2024) .
Medical-Specific VLMs. In contrast to general-purpose models, medical-specific VLMs are tailored for clinical applications, emphasizing domain adaptation and integration of specialized medical knowledge. Trained on curated medical datasets and aligned with diagnostic workflows, these models prioritize accuracy and reliability in healthcare settings. We select following medical-specific VLMs for evaluation, including LLaVA-Med (Li et al. 2023), Med-Flamingo (Moor et al. 2023), Qilin-Med-VL (Liu et al. 2023b), RadFM (Wu et al. 2023), MedDr (He et al. 2024), HuatuoGPT-Vision (Chen et al. 2024), MedVLM-R1 (Pan et al. 2025), and latest MedGemma series (Sellergren et al. 2025) and Lingshu series (Xu et al. 2025).
Closed-Source VLMs. Closed-source VLMs are developed and maintained by enterprises with inaccessible source code, typically provided to users via APIs for integration into applications. We select following closed-source VLMs for evaluation, including Gemini 2.5 Pro (Comanici et al. 2025), GPT-4o (Hurst et al. 2024), Moonshot-v1 (MoonshotAI 2025) and Qwen-VL-Max (Bai et al. 2023).